A 12 000 - kg airplane launched by a catapult from an aircraft carrier is accelerated from 0 to 200 km/h in 3 s. (a) How many times the acceleration of gravity is the airplane’s acceleration? (b) What is the average force that the catapult exerts on the air plane?
1
Expert's answer
2018-04-27T14:20:08-0400
(a)
a=(v-v_0)/t=(200-0)/3 1/3.6=18.52 m/s^2
a/g=18.52/9.81=1.9
(b) The average force that the catapult exerts on the air plane is
F=ma=(12000)(18.52)=220000 N=220 kN.